OpenStudy (anonymous):

(-2x^2+x+6)/(x-2) y int:-3 zeroes:-3/2 (I think) Vertical Asymptotes: None Holes:x=2 Horizontal/oblique Asymptotes: y=0 ( I think)

OpenStudy (anonymous):

can anyone tell me how the graph will look like its kinda confusing or check my answers at the very least

OpenStudy (anonymous):

\[\frac{-2x^2 + x +6}{x-2} = -2x -3\] \[2x = -3\] \[x = \frac{-3}{2}\]

OpenStudy (anonymous):

ah Sure

OpenStudy (anonymous):

A straight Line y = -2x - 3

OpenStudy (anonymous):

no sorry i wanna know since it has x intercept and aymptote at y=0

OpenStudy (anonymous):

Intercepts Y: (0.-3) X:(\(\frac{-3}{2}\),0)